# Privacy-Preserving Video Anomaly Detection: A Survey

### v1

@@ VII-B2 Edge-Cloud Collaboration P2VAD Systems
Edge-cloud collaboration leverages the cloud’s powerful computational capabilities while performing local data processing at the edge, allowing different layers of computing units to cooperate for maximizing detection benefits and resource utilization. This approach can offer scalable VAD solutions. Additionally, only localized devices have direct access to raw data, enhancing system transparency and trustworthiness. Specifically, privacy-preserving algorithms execute on edge devices, and only encrypted or abstracted information is transmitted to the cloud for further analysis and model updates, reducing the risk of exposing sensitive video data while handling more complex large-scale anomaly detection tasks. As adaptive learning systems [158], especially online and continual learning techniques, continue to evolve, P2VAD systems are expected to improve performance in dynamic environments. These systems can continually update models based on new data, allowing them to adapt to evolving privacy threats and anomalous patterns. This adaptability is particularly crucial in smart city applications where environmental conditions and data distributions frequently change.

### v2

@@ VII-B2 Edge-Cloud Collaboration P2VAD Systems
Edge-cloud collaboration leverages the cloud’s powerful computational capabilities while performing local data processing at the edge, allowing different layers of computing units to cooperate for maximizing detection benefits and resource utilization. This approach can offer scalable VAD solutions [163]. Additionally, only localized devices have direct access to raw data, enhancing system transparency and trustworthiness. Specifically, privacy-preserving algorithms execute on edge devices, and only encrypted or abstracted information is transmitted to the cloud for further analysis and model updates, reducing the risk of exposing sensitive video data while handling more complex large-scale anomaly detection tasks. As adaptive learning systems, especially online and continual learning techniques, continue to evolve, P2VAD systems are expected to improve performance in dynamic environments. These systems can continually update models based on new data, allowing them to adapt to evolving privacy threats and anomalous patterns.